OpenStudy (anonymous):

The inclination of l1 is arctan (1/3).if l2 makes an angle of 45 degree with l1, find the slope of l2.. with medal

OpenStudy (anonymous):



OpenStudy (anonymous):

the slope of L2 is the total inclination of L1 and 45 degress... the angle of L1 is... \[\theta = \tan^{-1} (1/3)=18.435\deg\] let say slope of line 2 is tan alpha... \[\alpha = 18.435+45=63.435\deg\] \[m _{L2}=\tan \alpha=\tan 63.435\deg=1.999999999999983\approx2.0\]